Please help me with this answer


♠️=7

♦️=5+♠️

♠️=1+♣️♣️♣️

♠️+♦️+♣️=?


1
Expert's answer
2022-04-05T15:43:35-0400

"\spadesuit =7\\\diamondsuit =5+\spadesuit \Rightarrow \diamondsuit =5+7=12\\\spadesuit =1+\clubsuit \clubsuit \clubsuit \Rightarrow \clubsuit \clubsuit \clubsuit =\spadesuit -1=6\Rightarrow \clubsuit =\frac{6}{3}=2\\\spadesuit +\diamondsuit +\clubsuit =7+2+12=21"